What is the proportion of variance for each of the following values in a 2 x 2 chi-square test for independence?

(a) X2 =3.96, N=50

(b) X 2 =5.23, N=75

(c) X2 = 12.00, N=100
Answer #1

Given

2 X 2 chi square

D.F = (2-1) (2-1) = 1

a)

X2 =3.96, N=50

Proportion of variance, V = sqrt (X2 / n*df) = sqrt ( 3.96 / 50*1) = 0.2814

b)

X 2 =5.23, N=75

Proportion of variance, V = sqrt ( 5.23 / 75*1) = 0.2641

c)

X2 = 12.00, N=100

Proportion of variance, V = sqrt ( 12/ 100*1) = 0.3464